OpenStudy (anonymous):

5x2+7x-3=0 wats te discriminant

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

so how do u find if it is a real or imaginary

OpenStudy (anonymous):

\[ ax^2+bx+c\]\[Discriminant = \,b^2-4ac. \]

OpenStudy (anonymous):

so how do i find it with tthat only and how is it real or imaginary

OpenStudy (campbell_st):

well in your question a = 5, b = 7 and c = -3 substitute and evaluate

OpenStudy (anonymous):

oh so the a will be 5

OpenStudy (anonymous):

put u dont have to take out the five and put it into a diffrent equation

OpenStudy (campbell_st):

not quite... I think you need to try again...

OpenStudy (anonymous):

oh ok thank u that was really helpful bro

OpenStudy (anonymous):

wait so whats real and imaginary mean

OpenStudy (campbell_st):

ok ithe calculation is \[7^2 - 4 \times 5 \times (-3) =\] if its greater than zero it has real roots if its equal to zero it has 1 repeated root is it less than zero then you complex roots...

OpenStudy (anonymous):

oh i get it thanks u mind when i need help i can ask u bro

OpenStudy (campbell_st):

thats ok... if I can do it

OpenStudy (anonymous):

lmao thanks bro

OpenStudy (anonymous):

i have another quetsion

OpenStudy (campbell_st):

ok

OpenStudy (anonymous):

ok i solved another question and i got 0 but it says to write if it is a real number tell weather they will be ration or irrational how would i do that im confused

OpenStudy (campbell_st):

ok.... so if you get a zero value the quadratic is a perfect square... whats the equation...?

OpenStudy (anonymous):

so would that be a real number so is it rational or irrational

OpenStudy (campbell_st):

here is an example y=x^2 + 6x + 9 or y = (x + 3)^2 has a repeated root at x = -3 and the discriminant is 36 - 36 = 0

OpenStudy (anonymous):

ok so is it rational or irrational

OpenStudy (campbell_st):

so a parabola such as \[y=x^2 - 6x + 5 ....= (x -5)(x -1) \] has roots at 1, 5 the discriminant 36 - 20 = 16 16 is a perfect square so the parabola has 2 unequal roots that are real and rational.... if you think about your original question the discriminant was 109.... which means 2 unequal real roots which are irrational as 109 isn't a square number.

OpenStudy (anonymous):

oh ok so an unperfect square is a irrational and a perfect square would be rational

OpenStudy (anonymous):

ok so x2 would be a=1 right

OpenStudy (campbell_st):

yes... as long as its >0

OpenStudy (anonymous):

ok y that face

OpenStudy (campbell_st):

good luck

OpenStudy (anonymous):

lmao thanks ur great help bro